# Determination of the stably free cancellation property for orders

### v1

@@ 11.3. Group structure of unit groups of finite rings
Let $\overline{\Lambda}=\Lambda/\mathfrak{h}$. Then $\overline{\Lambda}$ is a finite ring and an important challenge to compute its units $\overline{\Lambda}^{\times}$ as an abstract group. The point is that, although we can compute generators of $\overline{\Lambda}^{\times}$ by Corollary 11.4, we require a representation of $\overline{\Lambda}^{\times}$ that will allow us to efficiently determine the group-theoretic properties required for applications, such as the computation of the abelianisation $\overline{\Lambda}^{\times\mathrm{ab}}$. Specifically, representations which are useful for actual computations include realisations of $\overline{\Lambda}^{\times}$ as a subgroup of either $\mathrm{GL}_{n}(\mathbb{F}_{q})$ for some finite field $\mathbb{F}_{q}$, or of symmetric groups $\mathfrak{S}_{d}$, both of course with reasonable sizes for $n$ and $\lvert k\rvert$, or $d$, respectively (see [HEO05]).
@@
In general, the method of (i) allows computations to be performed more efficiently than the method of (ii), but of course the method of (i) is not always possible. We also note that if $\overline{\Lambda}$ can be written as a direct product of rings $R_{1}\times\cdots\times R_{s}$, then
@@
12. Improved freeness testing for generalised Swan modules
@@ 12.1. Setup
Let $K$ be a number field with ring of integers $\mathcal{O}=\mathcal{O}_{K}$, and let $A$ be a finite-dimensional semisimple $K$-algebra. Let $\Lambda$ be an $\mathcal{O}$-order in $A$ and let $\mathcal{M}$ be a maximal $\mathcal{O}$-order in $A$ containing $\Lambda$. Recall from Definition 8.1 that for $\beta\in\mathcal{M}\cap A^{\times}$, the generalised Swan module $X_{\beta}=X_{\beta,\Lambda,\mathcal{M}}$ is defined to be $\mathcal{M}\beta\cap\Lambda$.
@@ 12.2. Criteria for certain ideals to be free
We shall consider left ideals $X$ of $\Lambda$ with the following two properties: $X+\mathfrak{f}=\Lambda$ and there exists $\beta\in\mathcal{M}\cap A^{\times}$ such that $\mathcal{M}X=\mathcal{M}\beta$. Note that by Proposition 8.3, the generalised Swan modules $X_{\beta}$ considered in Algorithms 8.9, 9.1 and 10.2 always satisfy both these properties.
@@
Let $X$ be a left ideal of $\Lambda$. Suppose that $X+\mathfrak{f}=\Lambda$ and that there exists $\beta\in\mathcal{M}\cap A^{\times}$ such that $\mathcal{M}X=\mathcal{M}\beta$. Write $\beta=(\beta_{1},\beta_{2})\in\mathcal{M}_{1}\oplus\mathcal{M}_{2}$. Then the following are equivalent:
@@
    $X$ is free over $\Lambda$,
@@
    there exists $\overline{\alpha}\in(\Lambda/\mathfrak{f})^{\times}$ such that $\overline{\beta}\overline{\alpha}\in T$,
@@
    $(\overline{\alpha}_{1},\overline{\alpha}_{2})\in(\Lambda/\mathfrak{f})^{\times}\text{ and }\overline{\beta}_{1}\overline{\alpha}_{1}\in T_{1}.$
@@
The equivalence of (i) and (ii) follows from Proposition 7.1. Statement (iii) is just an immediate reformulation of (ii). ∎
[BJ11, Algorithm 3.1] essentially uses the the equivalence of conditions (i) and (ii) in Lemma 12.1 to reduce the freeness test to an enumeration over the set $T$ of cardinality $\lvert T\rvert=\lvert T_{1}\rvert\cdot\lvert T_{2}\rvert$. In the following, we show how condition (iii) can be used to replace this by an enumeration over $T_{2}$ and computations related to the reduced norm.
@@
which is surjective by [Bas68, Chapter III, (2.9) Corollary]. Let $\Sigma=\ker(\overline{\pi}_{2})$ and let
@@
Let $X$ be a left ideal of $\Lambda$. Suppose that $X+\mathfrak{f}=\Lambda$ and that there exists $\beta\in\mathcal{M}\cap A^{\times}$ such that $\mathcal{M}X=\mathcal{M}\beta$. Then $\overline{\beta}\in\left(\mathcal{M}/\mathfrak{f}\right)^{\times}$. Write $\beta=(\beta_{1},\beta_{2})\in\mathcal{M}_{1}\oplus\mathcal{M}_{2}$. Let $\overline{\alpha}_{2}\in\overline{\beta}_{2}^{-1}T_{2}$ and assume that $\overline{\gamma}_{1}\in\left(\mathcal{M}_{1}/\mathfrak{f}_{1}\right)^{\times}$ satisfies $(\overline{\gamma}_{1},\overline{\alpha}_{2})\in\left(\Lambda/\mathfrak{f}\right)^{\times}$. Then the following are equivalent:
@@
    There exists $\overline{\alpha}_{1}\in\left(\mathcal{M}_{1}/\mathfrak{f}_{1}\right)^{\times}$ such that $(\overline{\alpha}_{1},\overline{\alpha}_{2})\in(\Lambda/\mathfrak{f})^{\times}\text{ and }\overline{\beta}_{1}\overline{\alpha}_{1}\in T_{1}$,
@@
Assume now that (ii) holds. Then there exists a unit $\epsilon_{1}\in\mathcal{O}_{C_{1}}^{\times}$ and an element $(\overline{\delta}_{1},\overline{1})\in\left(\Lambda/\mathfrak{f}\right)^{\times}$ such that $\overline{\mathrm{nr}}(\overline{\beta}_{1}\overline{\gamma}_{1}\overline{\delta}_{1})=\overline{\epsilon}_{1}$ in $\left(\mathcal{O}_{C_{1}}/\mathfrak{g}_{1}\right)^{\times}$. By Eichler’s Theorem on units [CR87, (51.34)] there exists $u_{1}\in\mathcal{M}_{1}^{\times}$ such that $\overline{\beta}_{1}\overline{\gamma}_{1}\overline{\delta}_{1}=\overline{u}_{1}$, that is, $\overline{\beta}_{1}\overline{\gamma}_{1}\overline{\delta}_{1}\in T_{1}$. Since $(\overline{\gamma}_{1}\overline{\delta}_{1},\overline{\alpha}_{2})=(\overline{\gamma}_{1},\overline{\alpha}_{2})(\overline{\delta}_{1},\overline{1})\in\left(\Lambda/\mathfrak{f}\right)^{\times}$ we may take $\overline{\alpha}_{1}:=\overline{\gamma}_{1}\overline{\delta}_{1}$.
@@
Let $X$ be a left ideal of $\Lambda$. Suppose that $X+\mathfrak{f}=\Lambda$ and that there exists $\beta\in\mathcal{M}\cap A^{\times}$ such that $\mathcal{M}X=\mathcal{M}\beta$. Then the following statements hold.
@@
    $X$ is free over $\Lambda$ if and only if there exist $\overline{\alpha}_{2}\in\overline{\beta}_{2}^{-1}T_{2}$ and $\overline{\gamma}_{1}\in\left(\mathcal{M}_{1}/\mathfrak{f}_{1}\right)^{\times}$ such that $(\overline{\gamma}_{1},\overline{\alpha}_{2})\in\left(\Lambda/\mathfrak{f}\right)^{\times}$ and $\overline{\mathrm{nr}}(\overline{\beta}_{1}\overline{\gamma}_{1})\in\mathrm{im}(\sigma)$.
@@
    Fix $\overline{\alpha}_{2}\in\overline{\beta}_{2}^{-1}T_{2}$. Suppose there exists $\overline{\gamma}_{1}\in\left(\mathcal{M}_{1}/\mathfrak{f}_{1}\right)^{\times}$ such that $(\overline{\gamma}_{1},\overline{\alpha}_{2})\in\left(\Lambda/\mathfrak{f}\right)^{\times}$ and $\overline{\mathrm{nr}}(\overline{\beta}_{1}\overline{\gamma}_{1})\in\mathrm{im}(\sigma)$. Then for every $\overline{\mu}_{1}\in\left(\mathcal{M}_{1}/\mathfrak{f}_{1}\right)^{\times}$ such that $(\overline{\mu}_{1},\overline{\alpha}_{2})\in\left(\Lambda/\mathfrak{f}\right)^{\times}$, we have $\overline{\mathrm{nr}}(\overline{\beta}_{1}\overline{\mu}_{1})\in\mathrm{im}(\sigma)$.
@@
(i) Suppose that $X$ is free over $\Lambda$. Then by Lemma 12.1 there exist $\overline{\alpha}_{2}\in\overline{\beta}_{2}^{-1}T_{2}$ and $\overline{\alpha}_{1}\in(\mathcal{M}_{1}/\mathfrak{f}_{1})^{\times}$ such that $(\overline{\alpha}_{1},\overline{\alpha}_{2})\in(\Lambda/\mathfrak{f})^{\times}\text{ and }\overline{\beta}_{1}\overline{\alpha}_{1}\in T_{1}$. Moreover, $\overline{\mathrm{nr}}(\overline{\beta}_{1}\overline{\alpha}_{1})\bmod S\in\mathrm{im}(\sigma)$ by Lemma 12.2, and so we may take $\overline{\gamma}_{1}=\overline{\alpha}_{1}$. The converse holds again by Lemma 12.2.
@@
Let $\alpha_{2}\in\mathcal{M}_{2}$ with $\overline{\alpha}_{2}\in\left(\mathcal{M}_{2}/\mathfrak{f}\right)^{\times}$. The following are equivalent:
@@
    There exists an element $\overline{\gamma}_{1}\in(\mathcal{M}_{1}/\mathfrak{f}_{1})^{\times}$ such that $(\overline{\gamma}_{1},\overline{\alpha}_{2})\in\left(\Lambda/\mathfrak{f}\right)^{\times}$,
@@
Suppose that $A_{1}$ is Eichler and that every simple component of $A_{2}$ is a totally definite quaternion algebra. Let $X$ be a left ideal of $\Lambda$. Suppose that $X+\mathfrak{f}=\Lambda$ and that there exists $\beta\in\mathcal{M}\cap A^{\times}$ such that $\mathcal{M}X=\mathcal{M}\beta$. Write $\beta=(\beta_{1},\beta_{2})\in\mathcal{M}_{1}\oplus\mathcal{M}_{2}$. If $X$ is free over $\Lambda$ then the following algorithm returns $\mathsf{true}$; otherwise it returns $\mathsf{false}$.
@@
Compared to [BJ11, Algorithm 3.1] where the size of the enumeration set is essentially the cardinality of the image of the canonical map ×$\mathcal{M}^{\times}\longrightarrow\left(\mathcal{M}/\mathfrak{f}\right)^{\times}$, that is, $\lvert T_{1}\rvert\cdot\lvert T_{2}\rvert$, the enumeration set in Algorithm 12.12 is of size $\lvert T_{2}\rvert$, and hence only depends on the component $A_{2}$. Note that $T_{2}$ can be computed as explained in [BJ11, §4.5] (see also [KV10, Remark 7.5]), and this is much faster than computing the image of $\mathcal{M}\longrightarrow(\mathcal{M}/\mathfrak{f})^{\times}$ for $\mathcal{M}$ a maximal $\mathcal{O}$-order in an arbitrary finite-dimensional semisimple $K$-algebra.
@@ 12.3. Computation of $S$
There exist coprime primary ideals $\{\mathfrak{q}_{1},\dotsc,\mathfrak{q}_{r+s}\}$ and $\{\tilde{\mathfrak{q}}_{1},\dotsc,\tilde{\mathfrak{q}}_{r}\}$ of $\Lambda\cap C$ such that the following hold:
@@
    we have $\mathfrak{q}_{i}\subseteq\tilde{\mathfrak{q}}_{i}$ for $1\leq i\leq r$,
@@
    $\Lambda/\mathfrak{f}\cong\prod_{i=1}^{r+s}\Lambda/(\mathfrak{q}_{i}\Lambda+\mathfrak{f})\quad\text{and}\quad\Lambda/\tilde{\mathfrak{f}}\cong\prod_{i=1}^{r}\Lambda/(\tilde{\mathfrak{q}}_{i}\Lambda+\tilde{\mathfrak{f}}),$
@@
Consider $\mathfrak{g}=\mathfrak{f}\cap C$ and $\tilde{\mathfrak{g}}=\mathfrak{\tilde{f}}\cap C$, which are full $(\Lambda\cap C)$-ideals with $\mathfrak{g}\subseteq\tilde{\mathfrak{g}}$. Let $\mathfrak{p}_{1},\dotsc,\mathfrak{p}_{r}$ and $\mathfrak{p}_{1},\dotsc,\mathfrak{p}_{r+s}$ be the maximal ideals of $\Lambda\cap C$ containing $\tilde{\mathfrak{g}}$ and $\mathfrak{g}$ respectively. Let now $\mathfrak{g}=\prod_{i=1}^{r+s}\mathfrak{q_{i}}$ be a primary decomposition, where $\mathfrak{q}_{i}$ is $\mathfrak{p}_{i}$-primary for $1\leq i\leq r+s$. Since the $\mathfrak{p}_{i}$ are maximal, it is clear that for each $1\leq i\leq r$ also $\mathfrak{q}_{i}+\mathfrak{\tilde{g}}$ is $\mathfrak{p}_{i}$-primary. Thus $\tilde{\mathfrak{g}}=\mathfrak{g}+\tilde{\mathfrak{g}}=\prod_{i=1}^{r+s}\mathfrak{q}_{i}+\tilde{\mathfrak{g}}=\prod_{i=1}^{r}\tilde{\mathfrak{q}}_{i}$ is a primary decomposition of $\tilde{\mathfrak{g}}$, where $\tilde{\mathfrak{q_{i}}}:=\mathfrak{q}_{i}+\tilde{\mathfrak{g}}$ is $\mathfrak{p}_{i}$-primary for $1\leq i\leq r$. This immediately shows (i). For (ii), note that [BB06, 3.5 Lemma] implies that
@@
thus the claim follows from the Chinese remainder theorem. Finally, (iii) follows from (ii) together with Lemma 12.7. ∎
Since we know how to compute generators for each of the groups $\left(\Lambda/(\mathfrak{q}_{r+i}\Lambda+\mathfrak{f})\right)^{\times}$ by the results of §11.2, it remains to consider the computation of $\ker(\tau_{i})$ for $i=1,\ldots,r$. Each element in $\ker(\tau_{i})$ is the image in $(\Lambda/(\mathfrak{q}_{i}\Lambda+\mathfrak{f}))^{\times}$ of an element of the form $1+\lambda$ with $\lambda\in(\tilde{\mathfrak{q}}_{i}\Lambda+\tilde{\mathfrak{f}})$. Note, however, that not every element in $1+(\tilde{\mathfrak{q}}_{i}\Lambda+\tilde{\mathfrak{f}})$ is actually a unit mod $(\mathfrak{q}_{i}\Lambda+\mathfrak{f})$.
@@
Let $\mathfrak{p}$ be a maximal ideal of $\Lambda\cap C$, let $\mathfrak{q}$ and $\tilde{\mathfrak{q}}$ be two $\mathfrak{p}$-primary ideals of $\Lambda\cap C$ with ${\mathfrak{q}}\subseteq\tilde{\mathfrak{q}}$ and let
@@
the canonical projection. Furthermore set
@@
Then the following hold:
@@
(i) Since both $\tilde{\mathfrak{q}}$ and $\mathfrak{q}$ are $\mathfrak{p}$-primary, there exists $m>0$ such that $\tilde{\mathfrak{q}}^{m}\subseteq\mathfrak{q}$. Hence $\mathfrak{Q}_{1}^{m}\subseteq\mathfrak{Q}$ and every element of $\mathfrak{Q}_{1}/\mathfrak{Q}$ is nilpotent in $\Lambda/\mathfrak{Q}$. If $\lambda$ is invertible mod $\mathfrak{Q}_{1}$, then there exist $\mu\in\Lambda$ and $\xi\in\mathfrak{Q}_{1}$ such that $\lambda\mu=1-\xi$. Thus in $\Lambda/\mathfrak{Q}$ we have $\overline{\lambda}\overline{\mu}\cdot\sum_{n=0}^{\infty}\overline{\xi}^{n}=(\overline{1}-\overline{\xi})\cdot\sum_{n=0}^{\infty}\overline{\xi}^{n}=\overline{1}$. The converse is clear. (ii) now follows easily from (i). ∎

### v2

@@ 11.3. Group structure of unit groups of finite rings
Let $\overline{\Lambda}=\Lambda/\mathfrak{h}$. Then $\overline{\Lambda}$ is a finite ring and an important challenge is to compute its units $\overline{\Lambda}^{\times}$ as an abstract group. The point is that, although we can compute generators of $\overline{\Lambda}^{\times}$ by Corollary 11.4, we require a representation of $\overline{\Lambda}^{\times}$ that will allow us to efficiently determine the group-theoretic properties required for applications, such as the computation of the abelianisation $\overline{\Lambda}^{\times\mathrm{ab}}$. Specifically, representations which are useful for actual computations include realisations of $\overline{\Lambda}^{\times}$ as a subgroup of either $\mathrm{GL}_{n}(\mathbb{F}_{q})$ for some finite field $\mathbb{F}_{q}$, or of symmetric groups $\mathfrak{S}_{d}$, both of course with reasonable sizes for $n$ and $q$, or $d$, respectively (see [HEO05]).
@@
In general, the method of (i) allows computations to be performed more efficiently than the method of (ii), but of course the method of (i) is not always applicable. We also note that if $\overline{\Lambda}$ can be written as a direct product of rings $R_{1}\times\cdots\times R_{s}$, then
@@
12. Freeness testing using algebra decompositions
@@ 12.1. Setup
Let $K$ be a number field with ring of integers $\mathcal{O}=\mathcal{O}_{K}$, and let $A$ be a finite-dimensional semisimple $K$-algebra. Let $\Lambda$ be an $\mathcal{O}$-order in $A$ and let $\mathcal{M}$ be a maximal $\mathcal{O}$-order in $A$ containing $\Lambda$. Recall from Definition 8.1 that for $\beta\in\mathcal{M}\cap A^{\times}$, the test lattice $X_{\beta}=X_{\beta,\Lambda,\mathcal{M}}$ is defined to be $\mathcal{M}\beta\cap\Lambda$.
@@ 12.2. Criteria for certain ideals to be free
We shall consider left ideals $X$ of $\Lambda$ with the following two properties: $X+\mathfrak{f}=\Lambda$ and there exists $\beta\in\mathcal{M}\cap A^{\times}$ such that $\mathcal{M}X=\mathcal{M}\beta$. Note that by Proposition 8.3, the test lattices $X_{\beta}$ considered in Algorithms 8.9, 9.1 and 10.2 always satisfy both of these properties.
@@
Let $X$ be a left ideal of $\Lambda$. Suppose that $X+\mathfrak{f}=\Lambda$ and that there exists $\beta\in\mathcal{M}\cap A^{\times}$ such that $\mathcal{M}X=\mathcal{M}\beta$. Write $\beta=(\beta_{1},\beta_{2})\in\mathcal{M}_{1}\oplus\mathcal{M}_{2}$. Then the following assertions are equivalent:
@@
    $X$ is free over $\Lambda$.
@@
    There exists $\overline{\alpha}\in(\Lambda/\mathfrak{f})^{\times}$ such that $\overline{\beta}\overline{\alpha}\in T$.
@@
    $(\overline{\alpha}_{1},\overline{\alpha}_{2})\in(\Lambda/\mathfrak{f})^{\times}\quad\text{and}\quad\overline{\beta}_{1}\overline{\alpha}_{1}\in T_{1}.$
@@
The equivalence of (i) and (ii) follows from Proposition 7.1. Assertion (iii) is just an immediate reformulation of (ii). ∎
[BJ11, Algorithm 3.1] essentially uses the the equivalence of conditions (i) and (ii) in Lemma 12.1 to reduce the freeness test to an enumeration over the set $T$ of cardinality $\lvert T\rvert=\lvert T_{1}\rvert\cdot\lvert T_{2}\rvert$. In the following, we show how condition (iii) can be used to replace this by an enumeration over $T_{2}$ and computations related to the reduced norm. (See also Remark 12.6.)
@@
which is surjective by [Bas68, Chapter III, (2.9) Corollary]. Let $\Sigma:=\ker(\overline{\pi}_{2})$ and let
@@
Let $X$ be a left ideal of $\Lambda$. Suppose that $X+\mathfrak{f}=\Lambda$ and that there exists $\beta\in\mathcal{M}\cap A^{\times}$ such that $\mathcal{M}X=\mathcal{M}\beta$. Then $\overline{\beta}\in\left(\mathcal{M}/\mathfrak{f}\right)^{\times}$. Write $\beta=(\beta_{1},\beta_{2})\in\mathcal{M}_{1}\oplus\mathcal{M}_{2}$. Let $\overline{\alpha}_{2}\in\overline{\beta}_{2}^{-1}T_{2}$ and assume that $\overline{\gamma}_{1}\in\left(\mathcal{M}_{1}/\mathfrak{f}_{1}\right)^{\times}$ satisfies $(\overline{\gamma}_{1},\overline{\alpha}_{2})\in\left(\Lambda/\mathfrak{f}\right)^{\times}$. Then the following assertions are equivalent:
@@
    There exists $\overline{\alpha}_{1}\in\left(\mathcal{M}_{1}/\mathfrak{f}_{1}\right)^{\times}$ such that $(\overline{\alpha}_{1},\overline{\alpha}_{2})\in(\Lambda/\mathfrak{f})^{\times}\text{ and }\overline{\beta}_{1}\overline{\alpha}_{1}\in T_{1}$.
@@
Now assume that (ii) holds. Then there exists a unit $\epsilon_{1}\in\mathcal{O}_{C_{1}}^{\times}$ and an element $(\overline{\delta}_{1},\overline{1})\in\left(\Lambda/\mathfrak{f}\right)^{\times}$ with $\overline{\delta}_{1}\in\left(\mathcal{M}_{1}/\mathfrak{f}_{1}\right)^{\times}$ such that $\overline{\mathrm{nr}}(\overline{\beta}_{1}\overline{\gamma}_{1}\overline{\delta}_{1})=\overline{\epsilon}_{1}$ in $\left(\mathcal{O}_{C_{1}}/\mathfrak{g}_{1}\right)^{\times}$. By Eichler’s theorem on units [CR87, (51.34)] there exists $u_{1}\in\mathcal{M}_{1}^{\times}$ such that $\overline{\beta}_{1}\overline{\gamma}_{1}\overline{\delta}_{1}=\overline{u}_{1}$ in $\left(\mathcal{M}_{1}/\mathfrak{f}_{1}\right)^{\times}$, that is, $\overline{\beta}_{1}\overline{\gamma}_{1}\overline{\delta}_{1}\in T_{1}$. Since $(\overline{\gamma}_{1}\overline{\delta}_{1},\overline{\alpha}_{2})=(\overline{\gamma}_{1},\overline{\alpha}_{2})(\overline{\delta}_{1},\overline{1})\in\left(\Lambda/\mathfrak{f}\right)^{\times}$ we may take $\overline{\alpha}_{1}:=\overline{\gamma}_{1}\overline{\delta}_{1}$.
@@
Let $X$ be a left ideal of $\Lambda$. Suppose that $X+\mathfrak{f}=\Lambda$ and that there exists $\beta\in\mathcal{M}\cap A^{\times}$ such that $\mathcal{M}X=\mathcal{M}\beta$. Then the following statements hold:
@@
    $X$ is free over $\Lambda$ if and only if there exist $\overline{\alpha}_{2}\in\overline{\beta}_{2}^{-1}T_{2}$ and $\overline{\gamma}_{1}\in\left(\mathcal{M}_{1}/\mathfrak{f}_{1}\right)^{\times}$ such that $(\overline{\gamma}_{1},\overline{\alpha}_{2})\in\left(\Lambda/\mathfrak{f}\right)^{\times}$ and $\overline{\mathrm{nr}}(\overline{\beta}_{1}\overline{\gamma}_{1})\bmod{S}\in\mathrm{im}(\sigma)$.
@@
    Fix $\overline{\alpha}_{2}\in\overline{\beta}_{2}^{-1}T_{2}$. Suppose there exists $\overline{\gamma}_{1}\in\left(\mathcal{M}_{1}/\mathfrak{f}_{1}\right)^{\times}$ such that $(\overline{\gamma}_{1},\overline{\alpha}_{2})\in\left(\Lambda/\mathfrak{f}\right)^{\times}$ and $\overline{\mathrm{nr}}(\overline{\beta}_{1}\overline{\gamma}_{1})\bmod{S}\in\mathrm{im}(\sigma)$. Then for every $\overline{\mu}_{1}\in\left(\mathcal{M}_{1}/\mathfrak{f}_{1}\right)^{\times}$ such that $(\overline{\mu}_{1},\overline{\alpha}_{2})\in\left(\Lambda/\mathfrak{f}\right)^{\times}$, we have $\overline{\mathrm{nr}}(\overline{\beta}_{1}\overline{\mu}_{1})\in\mathrm{im}(\sigma)$.
@@
(i) Suppose that $X$ is free over $\Lambda$. Then by Lemma 12.1 there exist $\overline{\alpha}_{2}\in\overline{\beta}_{2}^{-1}T_{2}$ and $\overline{\alpha}_{1}\in(\mathcal{M}_{1}/\mathfrak{f}_{1})^{\times}$ such that $(\overline{\alpha}_{1},\overline{\alpha}_{2})\in(\Lambda/\mathfrak{f})^{\times}\text{ and }\overline{\beta}_{1}\overline{\alpha}_{1}\in T_{1}$. Moreover, $\overline{\mathrm{nr}}(\overline{\beta}_{1}\overline{\alpha}_{1})\bmod S\in\mathrm{im}(\sigma)$ by Lemma 12.2, and so we may take $\overline{\gamma}_{1}=\overline{\alpha}_{1}$. The converse holds by Lemmas 12.1 and 12.2.
@@
Let $\alpha_{2}\in\mathcal{M}_{2}$ with $\overline{\alpha}_{2}\in\left(\mathcal{M}_{2}/\mathfrak{f}\right)^{\times}$. The following assertions are equivalent:
@@
    There exists an element $\overline{\gamma}_{1}\in(\mathcal{M}_{1}/\mathfrak{f}_{1})^{\times}$ such that $(\overline{\gamma}_{1},\overline{\alpha}_{2})\in\left(\Lambda/\mathfrak{f}\right)^{\times}$.
@@
Suppose that $A_{1}$ satisfies the Eichler condition and that every simple component of $A_{2}$ is a totally definite quaternion algebra. Let $X$ be a left ideal of $\Lambda$. Suppose that $X+\mathfrak{f}=\Lambda$ and that there exists $\beta\in\mathcal{M}\cap A^{\times}$ such that $\mathcal{M}X=\mathcal{M}\beta$. Write $\beta=(\beta_{1},\beta_{2})\in\mathcal{M}_{1}\oplus\mathcal{M}_{2}$. If $X$ is free over $\Lambda$ then the following algorithm returns $\mathsf{true}$; otherwise it returns $\mathsf{false}$.
@@
Compared to [BJ11, Algorithm 3.1] where the size of the enumeration set is essentially the cardinality of the image of the canonical map ×$\mathcal{M}^{\times}\longrightarrow\left(\mathcal{M}/\mathfrak{f}\right)^{\times}$, that is, $\lvert T_{1}\rvert\cdot\lvert T_{2}\rvert$, the enumeration set in Algorithm 12.5 is of size $\lvert T_{2}\rvert$, and hence only depends on the component $A_{2}$. Note that $T_{2}$ can be computed as explained in [BJ11, §4.5] (see also [KV10, Remark 7.5]), and this is much faster than computing the image of $\mathcal{M}\longrightarrow(\mathcal{M}/\mathfrak{f})^{\times}$ for $\mathcal{M}$ a maximal $\mathcal{O}$-order in an arbitrary finite-dimensional semisimple $K$-algebra. If $|T_{1}|$ is ‘small’, then it may be that [BJ11, Algorithm 3.1] is faster than Algorithm 12.5 thanks to the optimisation described in [BJ11, §7] (also see [HJ20, §9]).
@@ 12.3. Computation of $S$
There exist coprime primary ideals $\{\mathfrak{q}_{1},\dotsc,\mathfrak{q}_{r+s}\}$ and $\{\tilde{\mathfrak{q}}_{1},\dotsc,\tilde{\mathfrak{q}}_{r}\}$ of $\Lambda\cap C$ such that the following assertions hold:
@@
    We have $\mathfrak{q}_{i}\subseteq\tilde{\mathfrak{q}}_{i}$ for $1\leq i\leq r$.
@@
    $\Lambda/\mathfrak{f}\cong\prod_{i=1}^{r+s}\Lambda/(\mathfrak{q}_{i}\Lambda+\mathfrak{f})\quad\text{and}\quad\Lambda/\tilde{\mathfrak{f}}\cong\prod_{i=1}^{r}\Lambda/(\tilde{\mathfrak{q}}_{i}\Lambda+\tilde{\mathfrak{f}}).$
@@
Consider $\mathfrak{g}=\mathfrak{f}\cap C$ and $\tilde{\mathfrak{g}}=\mathfrak{\tilde{f}}\cap C$, which are full $(\Lambda\cap C)$-ideals with $\mathfrak{g}\subseteq\tilde{\mathfrak{g}}$. Let $\mathfrak{p}_{1},\dotsc,\mathfrak{p}_{r}$ and $\mathfrak{p}_{1},\dotsc,\mathfrak{p}_{r+s}$ be the maximal ideals of $\Lambda\cap C$ containing $\tilde{\mathfrak{g}}$ and $\mathfrak{g}$, respectively. Let now $\mathfrak{g}=\prod_{i=1}^{r+s}\mathfrak{q_{i}}$ be a primary decomposition, where $\mathfrak{q}_{i}$ is $\mathfrak{p}_{i}$-primary for $1\leq i\leq r+s$. Since the $\mathfrak{p}_{i}$ are maximal, it is clear from [AM69, Proposition (4.2)] that $\tilde{\mathfrak{q}}_{i}:=\mathfrak{q}_{i}+\mathfrak{\tilde{g}}$ is also $\mathfrak{p}_{i}$-primary for $1\leq i\leq r$. Thus
$\tilde{\mathfrak{g}}=\mathfrak{g}+\tilde{\mathfrak{g}}=\left(\prod_{i=1}^{r+s}\mathfrak{q}_{i}\right)+\tilde{\mathfrak{g}}=\prod_{i=1}^{r+s}\left(\mathfrak{q}_{i}+\tilde{\mathfrak{g}}\right)=\prod_{i=1}^{r}\tilde{\mathfrak{q}}_{i}$
is a primary decomposition of $\tilde{\mathfrak{g}}$. This immediately shows (i). For (ii), note that [BB06, Lemma 3.5] implies that
@@
and thus the claim follows from the Chinese remainder theorem. Finally, (iii) follows from (ii) together with Lemma 12.7. ∎
Since we know how to compute generators for each of the groups $\left(\Lambda/(\mathfrak{q}_{r+i}\Lambda+\mathfrak{f})\right)^{\times}$ for $i=1,\ldots,s$ by the results of §11.2, it remains to consider the computation of $\ker(\tau_{i})$ for $i=1,\ldots,r$. Each element in $\ker(\tau_{i})$ is the image in $(\Lambda/(\mathfrak{q}_{i}\Lambda+\mathfrak{f}))^{\times}$ of an element of the form $1+\lambda$ with $\lambda\in(\tilde{\mathfrak{q}}_{i}\Lambda+\tilde{\mathfrak{f}})$. Note, however, that not every element in $1+(\tilde{\mathfrak{q}}_{i}\Lambda+\tilde{\mathfrak{f}})$ is actually a unit mod $(\mathfrak{q}_{i}\Lambda+\mathfrak{f})$.
@@
Let $\mathfrak{p}$ be a maximal ideal of $\Lambda\cap C$, let $\mathfrak{q}$ and $\tilde{\mathfrak{q}}$ be two $\mathfrak{p}$-primary ideals of $\Lambda\cap C$ with ${\mathfrak{q}}\subseteq\tilde{\mathfrak{q}}$, and let
@@
be the canonical projection. Furthermore set
@@
Then the following assertions hold:
@@
(i) Since both $\tilde{\mathfrak{q}}$ and $\mathfrak{q}$ are $\mathfrak{p}$-primary, there exists $m>0$ such that $\tilde{\mathfrak{q}}^{m}\subseteq\mathfrak{q}$. Hence $\mathfrak{Q}_{1}^{m}\subseteq\mathfrak{Q}$ and every element of $\mathfrak{Q}_{1}/\mathfrak{Q}$ is nilpotent in $\Lambda/\mathfrak{Q}$. Suppose $\alpha$ is invertible mod $\mathfrak{Q}_{1}$. Then there exist $\beta\in\Lambda$ and $\xi\in\mathfrak{Q}_{1}$ such that $\alpha\beta=1-\xi$. For $x\in\Lambda$, let $\overline{x}$ denote its image in $\Lambda/\mathfrak{Q}$. Then we have
$\overline{\alpha}\overline{\beta}\cdot\sum_{n=0}^{\infty}\overline{\xi}^{n}=(\overline{1}-\overline{\xi})\cdot\sum_{n=0}^{\infty}\overline{\xi}^{n}=\overline{1}.$
The converse is clear. (ii) now follows easily from (i). ∎